# Frame Sampling Strategies Matter: A Benchmark for small vision language models

### v1

@@ Abstract
Comparing vision language models (VLMs) on videos is particularly complex, as the performances is jointly determined by the model’s visual representation capacity and the frame-sampling strategy used to construct the input. Current video benchmarks are suspected to suffer from substantial frame-sampling bias, as models are evaluated with different frame selection strategies. In this work, we propose the first frame-accurate benchmark of state-of-the-art small VLMs (SVLMs) for video question-answering, evaluated under controlled frame-sampling strategies. Our results confirm the suspected bias and highlight both data-specific and task-specific behaviors of SVLMs under different frame-sampling techniques. By open-sourcing our benchmarking code, we provide the community with a reproducible and unbiased protocol for evaluating video VLMs and emphasize the need for standardized frame-sampling strategies tailored to each benchmarking dataset in future research.
Index Terms—  vision-language models, frame-sampling bias, video question-answering, benchmark, frame-accurate evaluation
@@ 2.1.3 CSTA
We adapt the CSTA algorithm for video understanding. First, $n$ initial frames are uniformly sampled from the video. Each frame is then embedded with GoogLeNet [19], and the embeddings are fed to an attention module that produces an attention map over the concatenated frames. A mixing module combines these maps with the frame features and passes the result to a classifier, which assigns an importance score to each frame.
@@ 2.3 Inference setup
We evaluate four SVLMs (Qwen2.5-VL-3B, Ovis2-2B, InternVL3-2B, and SmolVLM2-2.2B) on two video benchmarks: VideoMME and MVBench. For each model, inference was performed on video frames obtained under four sampling strategies: Standard (uniform-FPS), standard (single frame), adaptive (MaxInfo), and adaptive (CSTA).
Under standard uniform-FPS sampling, we used $r=2$ (two frames are sampled each second) with $N_{\max}$ = 96 and $N_{\min}$ = 4. We denote this configuration as FPS:2:4:96 in the figures. For single-frame sampling, we select either the center frame or the first frame of the video.
For MaxInfo and CSTA, we set $n=1000$ as the initial pool of uniformly sampled frames, then applied the respective algorithm to select an informative subset. The final selection was capped at $N_{\max}=96$ frames, consistent with the setting used in uniform-FPS sampling.
The choice of $N_{\max}=96$ across all experiments is determined by the visual token capacity of the smallest model in our study (SmolVLM). Fixing this upper bound ensures that all models operate under a uniform token budget, thereby enabling fair cross-model comparisons. All inferences were run on a relatively low compute budget equivalent to approximatively 2 weeks of 4 Nvidia A10G GPUs.
@@ 3 Results
Table 1 summarizes SVLMs performances under several frame sampling strategies on the VideoMME and MVBench benchmarks. On VideoMME, uniform-FPS sampling is best for every model. Qwen2.5 achieved the highest accuracy at 60.9%, and all SVLMs benefit from evenly spaced frames. By contrast, on MVBench, the optimal strategy is model-specific. Qwen2.5 performs best with MaxInfo (69.0%). SmolVLM2 performs best with a single center frame (57.0%). InternVL3 peaks with CSTA (79.2%) and Ovis2 prefers uniform-FPS (79.7%).
@@
Figure 1 compares our uniform-FPS results with literature reported scores for these SVLMs on VideoMME and MVBench. We can observe that once sampling strategy is standardized, performance can shift in either direction, indicating strong sensitivity to frame selection. The largest shift is observed for Ovis2 on MVBench, where accuracy rises from about 65% to 80%. This gain is likely due to our uniform-FPS protocol providing more frames than the limited set used in the original report (12 frames maximum). In contrast, Qwen2.5 shows little difference, as our uniform-FPS setting is close to that reported in the literature (they used FPS:2:4:768, while we used FPS:2:4:96). Using a common token budget and a single sampling policy removes this confound and makes cross-model differences more meaningful.
Table 1: Comparison of VideoMME and MVBench results across frame sampling strategies. FPS is the uniform sampling strategy at 2 frames per second with a total number of frames ranging from 4 to 96 frames. Max Info and CSTA strategies take 1000 input frames and select up to 96 key frames.
Model VideoMME MVBench
@@
Qwen2.5 41.9 43.5 60.9 57.5 57.9 52.8 54.9 65.0 69.0 65.6
SmolVLM2 36.2 38.4 50.7 48.7 49.2 52.1 57.0 53.5 52.6 52.3
InternVL3 37.9 42.8 54.6 53.6 53.3 67.5 69.7 77.5 77.6 79.2
Ovis2 40.8 42.1 54.2 52.5 53.3 69.9 72.9 79.7 78.0 77.9
@@ 4 Discussion
We present a standardized benchmark of state-of-the-art SVLMs that controls frame-sampling bias under two protocols: standard sampling and adaptive sampling. On VideoMME, Qwen2.5 achieves the highest accuracy with uniform-FPS sampling. On MVBench, Ovis2 and InternVL3 both attain the best accuracy with Ovis2 under uniform-FPS, and InternVL3 under the CSTA adaptive strategy.
These differences plausibly reflect mismatches in model’s training data distributions and their alignment with each benchmark. MVBench emphasizes short, motion-rich clips and prompts that stress temporal reasoning, whereas Video-MME spans broader durations and a wider mix of skills. As a result, models whose pretraining favors brief, dynamic sequences may align better with MVBench, while those tuned for longer contexts may fare better on VideoMME. Model size may also contribute: although all models are small (4B parameters), Qwen2.5 (3.7B) is larger than Ovis2 and InternVL3 (2B each).
For sampling strategies, uniform FPS is the strongest approach on VideoMME across all evaluated models. In contrast, performance on MVBench is model dependent. This pattern indicates that VideoMME rewards steady temporal coverage: evenly spaced frames preserve event order and reduce the risk of missing brief, repeated motions. On MVBench, performance is model dependent: some models, including Qwen2.5 and InternVL3, benefit from adaptive sampling. However, the gain comes with higher compute and time costs, and adaptive sampling is roughly three times slower than standard sampling because of preprocessing for frame selection. Single-frame sampling (center or first) underperforms all other strategies because it provides too little motion information and context for reliable responses.
Another important point is the comparison between literature-reported results and our controlled evaluation. Overall, controlling for frame-sampling bias is effective: it reveals discrepancies relative to reported scores. For example, on MVBench, the literature reports Ovis2 below Qwen2.5 (64.9% vs. 67.0%), partly because Qwen2.5 used uniform-FPS sampling (up to 768) while Ovis2 was evaluated with only a few frames (up to 12). Under our controlled benchmarking protocol, we observe the opposite ranking: Ovis2 performs better when both models receive the same input frames.
On VideoMME, Qwen2.5 performance scales with the frame budget: increasing $N_{\max}$ from 16 to 256 yields marked gains across sampling strategies, reflecting improved temporal coverage. Beyond 256 frames, accuracy plateaus or decreases slightly, likely due to redundancy and noise.

### v2

@@ Abstract
Comparing vision language models (VLMs) on videos is particularly complex, as the performances is jointly determined by the model’s visual representation capacity and the frame-sampling strategy used to construct the input. Current video benchmarks are suspected to suffer from substantial frame-sampling bias, as models are evaluated with different frame selection strategies. In this work, we propose the first controlled frame-sampling benchmark of state-of-the-art small VLMs (SVLMs) for video question-answering. Our results confirm the suspected bias and highlight both data-specific and task-specific behaviors of SVLMs under different frame-sampling techniques. By open-sourcing our benchmarking code, we provide the community with a reproducible and unbiased protocol for evaluating video VLMs and emphasize the need for standardized frame-sampling strategies tailored to each benchmarking dataset in future research.
Index Terms—  vision-language models, frame-sampling bias, video question-answering, benchmark, controlled frame-sampling benchmark
@@ 2.1.3 CSTA
We adapt the pretrained CSTA model for video understanding. First, $n$ initial frames are uniformly sampled from the video. Each frame is then embedded with GoogLeNet [17], and the embeddings are fed to an attention module that produces an attention map over the concatenated frames. A mixing module combines these maps with the frame features and passes the result to a classifier, which assigns an importance score to each frame.
@@ 2.3 Inference setup
We evaluate four small vision–language models, namely Qwen2.5-VL-3B, Ovis2-2B, InternVL3-2B, and SmolVLM2-2.2B, on two video understanding benchmarks: VideoMME and MVBench. For each model, inference is conducted on video frames obtained using four sampling strategies: single-frame sampling, uniform FPS sampling, and two adaptive methods : MaxInfo and CSTA.
All models are evaluated using the official prompts provided by each benchmark, ensuring consistent prompt control across experiments. We rely exclusively on fixed pretrained checkpoints and deterministic inference. As a result, there is no variability induced by random seeds, and the dominant source of uncertainty arises from the finite size of the evaluation datasets. Accordingly, we estimate standard errors using a binomial approximation that depends solely on the dataset size.
For uniform FPS sampling, we use a frame rate of $r=2$, corresponding to two frames per second, with a minimum of $N_{\min}=4$ and a maximum of $N_{\max}=96$ frames per video. This configuration is denoted as FPS:2:4:96. In the single-frame setting, we evaluate two variants by selecting either the center frame or the first frame of each video.
For the adaptive sampling strategies MaxInfo and CSTA, we first construct an initial pool of $n=1000$ uniformly sampled frames to balance temporal coverage and computational cost. This choice is particularly important for MaxInfo, as it involves SVD and MaxVol operations with computational complexity that scales approximately quadratically with the number of frames. Each method then selects an informative subset of frames from this pool, with the final selection capped at $N_{\max}=96$ frames, matching the upper bound used in the uniform FPS setting.
The choice of $N_{\max}=96$ is dictated by the visual token capacity of the smallest model considered in our study, namely SmolVLM. Fixing this upper limit across all experiments enforces a uniform token budget for all models, thereby enabling fair and controlled cross-model comparisons. All experiments were conducted under a modest computational budget, corresponding to approximately two weeks of runtime on four NVIDIA A10G GPUs.
@@ 3 Results
3.1 Performance Analysis
Table 1 summarizes SVLMs performances under several frame sampling strategies on the VideoMME and MVBench benchmarks. On VideoMME, uniform-FPS sampling is best for every model. Qwen2.5 achieved the highest accuracy at 60.9%, and all SVLMs benefit from evenly spaced frames. By contrast, on MVBench, the optimal strategy is model-specific. Qwen2.5 achieves the best performance with MaxInfo (68.4%). SmolVLM2 and Ovis2 perform best with uniform-FPS (51.0% and 65.8%, respectively). InternVL3 reaches its highest score with CSTA (65.8%).
@@ 3.1 Performance Analysis
Figure 1 compares our uniform-FPS results with literature reported scores for these SVLMs on VideoMME and MVBench. We can observe that once sampling strategy is standardized, performance can shift in either direction, indicating strong sensitivity to frame selection. The largest shift is observed for InternVL3 on MVBench, where accuracy dropped from about 70% to 65%. This drop is likely due to a distribution shift: InternVL3 is evaluated with a fixed number of uniformly spaced frames (16 frames), whereas our fixed 2 fps regime yields 4 to 96 frames, under-samples fast actions in short clips. In contrast, Qwen2.5 shows little difference, as our uniform-FPS setting is close to that reported in the literature (they used FPS:2:4:768, while we used FPS:2:4:96). Using a common token budget and a single sampling policy removes this confound and makes cross-model differences more meaningful.
Table 1: Comparison of VideoMME and MVBench results across frame sampling strategies. FPS (fps:2:4:96): uniform sampling at 2 fps. Max Info (maxinfo:1000:96) and CSTA (csta:1000:96): adaptive selection of up to 96 frames from 1000 inputs. Standard errors are reported using binomial approximation to estimate the uncertainty inherent to the finite evaluation sets.
Model VideoMME ($\pm$1.66%) MVBench ($\pm$0.76%)
@@
Qwen2.5 41.9 43.5 60.9 57.5 57.9 47.5 50.9 64.3 68.4 64.0
SmolVLM2 36.2 38.4 50.7 48.7 49.2 42.9 45.0 51.5 51.0 50.5
InternVL3 37.9 42.8 54.6 53.6 53.3 44.9 48.4 65.1 63.1 65.8
Ovis2 40.8 42.1 54.2 52.5 53.3 47.3 51.8 65.8 62.8 62.3
@@
3.2 Runtime Analysis
Table 2 presents the wall-clock inference time (in hours) for VideoMME and MVBench across different frame-sampling strategies, corresponding to the experiments whose results are reported in Table 1. Single-frame inference is computationally negligible for all models, remaining below 0.36 h in every configuration. This highlights the low overhead of per-video inference when minimal temporal context is used.
In contrast, uniform FPS sampling incurs a substantial increase in runtime, with inference times ranging from approximately 5–7 h on VideoMME and 1.6–2.4 h on MVBench. This increase reflects the processing of up to 96 frames per video, confirming the near-linear relationship between frame count and inference cost.
The most computationally demanding strategies are Max Info and CSTA. These methods require an expensive pre-selection stage, leading to inference times exceeding 21 h on VideoMME and 6–12 h on MVBench across all models. Importantly, Max Info and CSTA exhibit comparable runtimes, indicating that their computational burden is dominated by the initial dense frame processing.
We also report in Figure 4 the inference runtime, focusing on Qwen2.5 performance on VideoMME as $N_{\max}$ increases from 64 to 600. The results show that dynamic sampling methods can become particularly useful when the frame budget is large (e.g., $N_{\max}=600$). In this regime, the additional preprocessing cost incurred during frame preselection allows these methods to reduce the number of frames ultimately processed by the SVLM. In contrast, FPS directly processes up to 600 frames without any preselection, leading to similarly high inference times at large $N_{\max}$.
Table 2: Inference time comparison (in hours) across frame sampling strategies. FPS (fps:2:4:96): uniform sampling at 2 fps. Max Info (maxinfo:1000:96) and CSTA (csta:1000:96): adaptive selection of up to 96 frames from 1000 inputs.
Model VideoMME MVBench
Single Frame FPS Max Info CSTA Single Frame FPS Max Info CSTA
Qwen2.5 0.23 5.22 21.93 21.30 0.31 2.05 11.04 7.03
SmolVLM 0.33 5.41 21.30 22.06 0.25 1.67 9.97 6.49
InternVL 0.26 6.27 23.10 22.56 0.36 1.98 10.85 7.04
Ovis 0.21 6.90 23.49 23.10 0.34 2.38 11.91 7.67
Fig. 4: Impact of maximum frame limit on inference time
@@ 4 Discussion
We present a standardized benchmark of state-of-the-art SVLMs that controls frame-sampling bias under two protocols: standard sampling and adaptive sampling. On VideoMME, Qwen2.5 attains the highest accuracy with uniform-FPS, outperforming other SVLMs by at least 6%. On MVBench, Qwen2.5 also leads with MaxInfo, exceeding other SVLMs by at least 3%. Ovis2 and InternVL3 reach comparable accuracies, both outperforming SmolVLM2 across the two benchmarks.
These results highlight the importance of the model’s training data distribution and its alignment with each benchmark. MVBench emphasizes short, motion-rich clips and prompts that stress temporal reasoning, whereas Video-MME spans broader durations and a wider mix of skills. As a result, models whose pretraining favors brief, dynamic sequences may align better with MVBench, while those tuned for longer contexts may fare better on VideoMME. Model size may also contribute: although all models are small (up to 4B parameters), Qwen2.5 (3.7B) is larger than Ovis2 and InternVL3 (2B each).
For sampling strategies, uniform FPS is the strongest approach on VideoMME across all evaluated models. In contrast, performance on MVBench is model dependent. This pattern indicates that VideoMME rewards steady temporal coverage: evenly spaced frames preserve event order and reduce the risk of missing brief, repeated motions. On MVBench, performance is model dependent: some models, including Qwen2.5 and InternVL3, benefit from adaptive sampling. However, the gain comes with higher compute and time costs, and adaptive sampling is roughly three to five times slower than standard sampling because of preprocessing for frame selection. Single-frame sampling (center or first) underperforms all other strategies because it provides too little motion information and context for reliable responses.
Another important point is the comparison between literature-reported results and our controlled evaluation. Overall, controlling for frame-sampling bias is effective: it reveals discrepancies relative to reported scores. On MVBench, prior reports place Ovis2 below InternVL3 and Qwen2.5 (64.9% vs. 70.0% and 67.0%), partly because Qwen2.5 used uniform-FPS sampling (up to 768 frames) and InternVL3 evaluated 16 uniformly spaced frames, whereas Ovis2 was evaluated with only a few frames (up to 12). Under our controlled protocol, the ranking changes: when all models receive the same input frames, Ovis2 performs comparably (Ovis2 65.8% vs. InternVL3 65.0% and Qwen2.5 64.3%).
On VideoMME, Qwen2.5 performance scales with the frame budget: increasing $N_{\max}$ from 16 to 256 yields marked gains across sampling strategies, reflecting improved temporal coverage. Beyond 256 frames, accuracy plateaus or decreases slightly, likely due to redundancy and noise. We also find that dynamic sampling could becomes particularly useful when the frame budget is large, since the cost incurred during frame preselection allows these methods to reduce the number of frames ultimately processed by the SVLM, making the final inference time comparable to that of uniform sampling.